Find an equation of the line through the point (3, 5) that cuts off the least area from the first quadrant?

1
Expert's answer
2022-06-07T00:26:05-0400

The general equation of a straight line is "y=mx+b," where m is a slope, and b is the intercept on the y-axis.

Suppose the line passing through the points (3, 5) and (0, b).




Then we have:

"5=3m+b,"

"3m=5-b,"

"m=\frac{5-b}{3}."

We have found the slope, so the equation of the curve is:

"y=\frac{5-b}{3}x+b."


Now we can find the intercept on the x-axis (y=0):

"0=\frac{5-b}{3}x+b,"

"\frac{5-b}{3}x=-b,"

"x=\frac{-3b}{5-b}=\frac{3b}{b-5}."


The area of the right triangle is "S=\frac{1}{2}ab" , so we have

"S=\frac{1}{2}b\frac{3b}{b-5}=\frac{3b^2}{2(b-5)}."

To find its minimum we have to find the derivative:

"S'=\frac{6b\cdot 2(b-5)-3b^2\cdot 2}{4(b-5)^2}=\frac{12b^2-60b-6b^2}{4(b-5)^2}="

"=\frac{6b^2-60b}{4(b-5)^2}=\frac{6b(b-10)}{4(b-5)^2}."


"S'=0" when b=0 or b=10.




So b=10 produce the least area, and

"m=\frac{5-b}{3}=\frac{5-10}{3}=-\frac{5}{3}."


The equation of the line through the point (3, 5) that cuts off the least area from the first quadrant is

"y=-\frac{5}{3}x+10."


Answer: "y=-\frac{5}{3}x+10."